# SADA: Semantic adversarial unsupervised domain adaptation for Temporal Action Localization

### 20 Dec 2023 v1

@@ 4.2 Experimental results
| Ours (SADA) | 31.54 | 30.68 | 28.77 | 25.52 | 20.22 | 27.34 |
| Scenario | Model | mAP10% | mAP20% | mAP30% | mAP40% | mAP50% | Avg |
@@
In this section, we present the main experimental results evaluating appearance and acquisition shifts. All these experiments follow the standard transductive unsupervised DA protocol [44, 10], and report the mean average precision (mAP) at different intersection over union (IOU) thresholds ().
Appearance shift scenarios. In Tab 1 we present the results obtained in the 4 different scenarios that we designed to evaluate the performance of our method when facing different domain shifts induced by changes in the background information. In this table, we compare our method to Actionformer [62] and TriDet [49], the two best-performing methods on EpicKitchens100 dataset [14] for TAL at the time of this writing.
In the first two experiments where we use dark-counter kitchens to define the source domains, we observe a consistent relative improvement of SADA over the chosen baseline methods of up to . We observe a similar behavior for the 2 scenarios with white-counter kitchens acting as source, where our model improves the baselines with up to .
Acquisition shifts. Similarly, Tab. 2 shows the results for the two proposed scenarios evaluating acquisition domain shifts – e.g., camera and lighting conditions. In the first scenario where we define old videos as the source domain, we observe a considerable relative improvement of over TriDet [49]. This improvement becomes more prominent comparing it to ActionFormer which we improve by . The results indicate that the opposite setup presents a much more challenging scenario. In this case, our method attains a slightly more moderate yet consistent improvement of up to .
@@ 4.3 Ablation studies
Comparing to other domain adaptation methods. In Sec. 4.2 we showed the performance increase that SADA yields with respect to SOTA source-only methods on our 6 setups. This comparison does not include any domain-adaptation-based method as our proposal is the first one suitable for sparse temporal detection setups. Nevertheless, this raises the question of how other domain adaptation methods would perform if coupled with our chosen multi-resolution architecture? In Tab. 3 we conduct several relevant experiments on Scenarios 1 and 3, defining the source domain as new and old black-counter kitchens, respectively. We refer to the Supp. for the ablations on other benchmarking scenarios.
@@
| Semantic centroids [58] | 31.07 | 27.16 | 17.21 | 25.14 |
| SSTDA [8] | 27.16 | 24.32 | 16.30 | 22.59 |
@@
| Semantic centroids [58] | 30.88 | 27.44 | 18.25 | 25.52 |
| SSTDA [8] | 29.50 | 26.87 | 17.13 | 24.50 |
@@
For this, we benchmark the performance of our method with its source-only variant. We also include a direct comparison with adaptations of several domain-adaptation methods like DANN [18], semantic centroid alignment [58], and SSTDA [8], the closest work to our proposal.
As observed in Tab. 3, in the first considered scenario our method consistently yields a better performance than any of the evaluated domain adaptation methods. More specifically, we find that the first two alternatives yield only a marginal relative improvement over the source-only variant of up , whereas SSTDA [9] obtains a considerable decay of performance. For the latter, we hypothesize that this is due to the change of paradigm. Despite its original success, SSTDA was designed for frame-level prediction setups, inherently different from the task at hand. We observe a similar behavior in the second scenario, where DANN seems to work considerably better than the other baselines, but still well below our proposal, SADA.
Analysis of our loss. In this section we ablate over different variants of our proposed SADA loss (see Eq. 9). Concretely, in Tab. 4 we study the effect of class-wise distribution alignment (Eq. 7), global distribution alignment (DANN [18]) and semantic background alignment (Eq. 8).
In this regard, we highlight that the global adaptation seems to consistently improve upon aligning only background embeddings. This is because the latter yields only a partial alignment of the embeddings, not considering class anchors. Therefore, a rougher yet complete adaptation might seem beneficial.
We also observe that aligning local class-wise distributions has a considerable positive effect, especially in the first scenario. Its effect, however, considerably decreases when combined with a global alignment [18], as all the class anchors are then subject to the concurrent alignments of domain-level and class-wise distributions. This observation is also consistent with the performance decrease that we observe when combining global and background alignment, which again suggests that the concurrent alignment of background embeddings with two adaptation losses is harmful to performance.
Finally, we observe that we consistently obtain the best results when the local alignment loss is coupled with the complementary (non-overlapping) background loss – i.e., SADA loss – indicating that this semantic fine-grained, yet complete, alignment is the most desirable approach. We refer to the Supp. for a more complete class-wise analysis.
@@
Qualitative analysis. We complement our quantitative results with a qualitative study. For this, we depict in Fig. 4 a segment visualization of our proposed method versus the chosen baseline models. In this visualization, we can observe that the Actionformer [62] misses many of the segments in the shown video clip ignoring all the take actions and mistaking a close action for a take. Tridet [49] performs better but misses all the take actions in the first half of the video while worsening the boundary prediction of the last close. SADA improves Tridet by predicting the second take action and considerably improving the boundary of the last close action.
In Fig. 5 we also show the TSNE projections of the domain-invariant embeddings of SADA with respect to its source-only version. Given that our main contribution is an adversarial-based class-wise loss, we depict the TSNE plots of the 3 majority action classes (first 3 columns). Observe that the source only (top row) yields clearly unaligned distributions with scarce to no overlap in the projected space. Our method, in contrast, presents a considerable distribution mix improving the alignment of class-wise distributions across domains. Given the anchor-based nature of our method, we have numerous background anchors – i.e., not assigned to any true action label. As observed in the last column, these are also aligned by our method (resulting from Eq. 8) effectively aligning the entire data distributions but in a semantically sensitive way. We refer to the Supp. for the complete qualitative study.
5 Conclusions and limitations
In this work, we deal for the first time with Unsupervised Domain Adaptation on realistic Temporal Action Localization (TAL) scenarios. We propose a novel semantic adversarial loss that enables a more fine-grained distribution alignment compared to existing global-distribution-based approaches. Given the lack of suitable evaluation setups for this scenario, we propose a suite of 6 different benchmarks that provide a comprehensive assessment of the model performance across various domain shifts. These experiments indicate that our model yields a considerable improvement over state-of-the-art methods, which we support with extensive quantitative and qualitative results.
Limitations and future work. Despite the virtues of our proposed benchmarking scenarios, we leave as future work to study alternative cross-dataset setups including larger and different domain shifts, e.g. extreme viewpoint changes or motion. This would possibly require relabeling new datasets. We also plan on complementing these studies with long-tail scenarios, analyzing the effect of domain adaptation optimization in the presence of highly unbalanced data in terms of labels. We also leave as future work the analysis and possible benefits of the adaptation of our semantic adversarial loss to other domain adaptation problems, even to those not necessarily related to video tasks.
Supplementary Material
In this supplementary material, we first provide the implementation details needed to reproduce our work (see Sec. A). We then provide additional details of the description of Sec. 4.1 regarding our 6 proposed benchmarking scenarios (see Sec. B), which we complement with a visualization of the different scenarios (available in this link). In Sec. C we detail the necessary adaptations that we made on existing unsupervised domain adaptation (UDA) so that they could operate on our proposed setups. We then extend our qualitative analysis from Sec. 4.3, providing additional segment visualizations (see Sec. D.1 and this link) and an additional study of the effect of the alignment loss SADA on multiple resolution levels (see Sec. D.2). In Sec. E we provide empirical evidences that justify the choice of using learnable class embeddings. Find also in Sec. F an extension of the main-paper ablation from Sec. 4.2, where we present the experimental results for two additional scenarios. We complement these results with a more thorough study of the per-class performance of our model on multiple scenarios (see Sec. G). Finally, in Sec. H we ablate over the potential negative impact that the presence of negative anchors has on the performance of our model.
A Implementation details
Our implementation of SADA is based on the PyTorch [45] framework. SADA extracts video features using a frozen Slowfast [5] video backbone pre-trained on Kinetics [27]. This backbone is fed with raw videos from EpicKitchens100 [14] with a rate of FPS, a feature stride of , and a maximum length of ensured by either padding – for shorter videos – or random cropping. To ensure a fair comparison with [49], we also define the main feature extractor as a 6-level SGP feature pyramid [49] that follows a max pooling strategy and uses a downsampling rate of and an internal feature dimensionality of . We model the classification and localization heads as a sequence of three 1D-CNNs with a kernel size of . These heads are shared across embeddings of different resolution levels. Finally, we introduce level-wise domain discriminators designed as a multi-layer perceptron (MLP) of depth . The weights of the losses are tuned following a grid search strategy.
During training, the model first shuffles the videos of the source and target domain, respectively. It then feeds a batch of videos of each of the domains, resulting in a batch of videos per training iteration. Note that depending on the scenario under study, one of the domains may have more data than the other. For this reason, in each epoch we repeat the smaller domain until the other finishes, ensuring that all the data of each of the domains is leveraged at least once. Moreover, the model architecture is optimized for epochs using AdamW [39] with a learning rate of with cosine decay and a warm-up phase of epochs.
B Extended benchmark description
In this section we describe in more depth the 6 different benchmarking scenarios that we propose in Sec. 4.1. For this, in Sec. B.1 we first briefly describe EpicKitchens100 [14], the base of all our benchmarks. We then illustrate in Sec. B.2 and Sec. B.3 the 2 different domain shifts that we identify, which we complement with a more detailed description of the specific scenarios derived from each of them.
B.1 Description EpicKitchens100
EpicKitchens100 [14] is a large-scale ego-centric dataset that has recently gained a lot of attraction in the Temporal Action Localization community to test for challenging detection scenarios. This dataset extends the previous version of the dataset, namely EpicKitchens55 [13]. This dataset contains videos of 32 subjects performing daily cooking activities – e.g., wash, put, open, close – in different kitchen environments. This results in 100 hours of non-scripted videos containing 20M frames annotated with 90K action segments.
One of the key challenges of this dataset is the presence of 97 different actions, forming a very long tail action distribution. This inherent characteristic is however an issue that falls beyond the scope of this paper, given the important additional challenges that this poses in existing domain adaptation-based methods. For this reason, in all our experiments we consider only the 10 majority classes – i.e., take, put, wash, open, close, insert, turn-on, cut, turn-off and pour. As shown in Fig. A this pruning keeps most of the labeled action segments of the original dataset, concretely keeping up to of it (marked in blue in the histogram). Moreover, we find that this design decision does not diminish in a relevant way the challenging nature of this task. Proof of this is that the currently best-performing methods on this dataset – i.e., Actionformer [62] and Tridet [49]– still attained mAP metrics inferior to in all our proposed scenarios.
B.2 Studying acquisition shifts
The first domain shift that we identify in the EpicKitchens100 [14] is what we call the acquisition shift. This shift refers to the changes induced by differences in the acquisition conditions of the videos. In the case of EpicKitchens100 [14] this results from the extension of the original dataset EpicKitchens55 [13]. The original dataset, concretely, was formed by 55 hours of non-scripted videos, and nearly 40K action segments. Hence, we find that this presents a suitable setup to define one domain as the old videos – recorded in the original EpicKitchens55 [13] – and the other formed by the new videos – recorded for the extended version. As argued by [14] this results in several important domain gaps that are captured by these data splits:
Changes in the acquisition devices: The new videos that were recorded during the extension relied on a newer camera device. Importantly, this camera incorporates camera stabilization techniques. Fig. B visually depicts the improvement in the stability of new over old videos.
B.3 Studying appearance shifts
Another important domain shift that we are interested in capturing is one induced by changes in the background. In this regard, there are several possibilities for this, but we find that the vast majority allow only a vague intuition of the domain shift that is under study. For instance, [14] argues that in the extension, several kitchens might have some changes in the furniture or the order of the main kitchen utensils. But, which kitchens really contain these changes? Not understanding this issue in sufficient depth results in an obscure evaluation, which we try to avoid in this work.
For this reason, we identify a clear, understandable domain gap. This is the color of the kitchen counters. This is an essential part of the background information, and thus, we hypothesize that adapting to this factor is also critical to obtain high-performing models. In this regard, we split the data into three different domains: dark, white, and other types of kitchen counters. This provides a clear domain gap that we can visualize and thus, understand (see Fig. C). We refer interested readers to this link for video content that further depicts the differences between these domains.
All in all, we stress that all our proposed scenarios provide sufficiently even splits to design single-domain setups where we can ensure that there is no easy generalization of the target domain by means of the source training data (as argued in Sec. 4.1)
C Unsupervised domain adaptation baseline description
One of the important challenges that we faced is that to the best of our knowledge, no previous unsupervised domain adaptation (UDA) work is directly comparable with our proposed setups. Hence, we adapted several existing UDA methods to our chosen multi-resolution architecture. This permits to establish a fair comparison with our main contribution, the novel Semantic Adversarial loss (SADA). Concretely, we adapted DANN [17], a semantic centroid alignment method [58] and SSTDA [8].
With this, we define the final loss for the multi-resolution DANN:
where is a hyperparameter that controls the influence of each of the resolution levels on the overall adaptation loss.
Semantic Centroid alignment [58]: In the original work from [58], they propose to compute an online centroid estimate of each of the class embeddings. This keeps important resemblances with our approach as they attempt to reduce the discrepancy between embeddings across domains that belong to a given action class. For this, they also require classifying each of the anchor embeddings to their corresponding class labels. In the context of UDA, this is straightforward in the case of the source domain as we have their corresponding labels. Doing so on the target domain is more challenging as this domain is unlabelled. Consequently, we follow the same strategy presented in Eq. 11 and Eq. 12, classifying the target domain embeddings according to their pseudo-label. Hence, classifying them according to the highest confidence predicted class, if this is above a given threshold . This results in the source embeddings of class and resolution level , and the corresponding embeddings of the target domain .
SSTDA [8]: In the original work, SSTDA [8] proposes to couple what they called Local SSTDA with Global SSTDA. Local SSTDA refers to the standard frame-level alignment loss. To adapt this to our multi-resolution setup, we follow the same strategy as in our adaptation of DANN. Global SSTDA, in contrast, refers to a sequential domain prediction loss. This splits the original source and target domain into 2 clips of sequential frames. They then apply a domain attentive temporal pooling [8] to first weight the frames according to their entropy, and then pool them into a single compact clip embedding. The source and target clip embeddings are then randomly shuffled, and finally passed through a domain classifier that learns to predict the domain that each of the clips belong to.
To adapt this methodology to our multi-resolution setup, we apply this for every level independently. Concretely, we split into two segments the domain-invariant feature representations of domains and for a given level , yielding and . Then, following the original proposal, we apply domain attentive temporal pooling to compute segment-level representations and . Finally, we apply the sequential domain prediction on these segment embeddings of a given resolution level . For this, we shuffle them randomly into – e.g., – and train a level-wise domain classifier to predict the domain labels – e.g., .
D Additional qualitative results
In this section we extend our qualitative analysis from Sec. 4.3. To this end, we first provide in Sec. D.1 additional segment visualizations. Then, in Sec. D.2 we extend the analysis of the TSNE plots to study the effect of multi-resolution in the overall alignment of embeddings.
D.1 Segment visualization
In Fig. D we present a segment visualization of scenario 3 – i.e., using old dark-counter kitchens as the source domain, and the rest as the target. Observe that in this case, both ActionFormer [62] and Tridet [49] perform similarly. Concretely, they mainly miss one close action at the beginning, a put in the middle section of the video, and finally the last two actions open and take. In contrast, our model is able to correctly detect the previously missing close, and the final take. In short, reducing by half the number of undetected actions in the video.
D.2 Extended TSNE analysis: Studying the effect of multi resolution
In this section, we extend the analysis of Sec. 4.3 to ablate over the effect that our alignment loss has on different resolution levels of the model. For this, in Figs. E, F and G we present the corresponding TSNE plots of the three majority action classes on 3 resolution levels. Here level 0 is the shallowest level and subsequent levels progressively reduce the resolution by 2. Observe that the influence of the alignment loss follows a similar pattern in all the resolutions. Concretely, in the three presented, we observe a considerable mixing between the source and target distributions (red and blue, respectively). This contrasts with the very clear disentanglement of the representations of both domains in the source-only version, where these present little to no overlap. We highlight what seems to be the only exception which is resolution level 3 of class 1. This one shows little improvement in the mixing of the distributions compared to the source-only variant.
We also find that our overall improvement of the mixing of the distributions is also consistent when analyzing the background class embeddings. Observe in Fig. H that in all 3 studied resolution levels, SADA improves very considerably the alignment, pushing the feature space of both domains to be domain invariant. This emphasizes the positive influence of the background alignment term of our loss (see Eq. 8 of the main paper).
E Study of the class embedding

### 05 Apr 2024 v2

@@ 4.2 Experimental results
| Ours (source-only) | 30.03 | 28.70 | 26.62 | 23.03 | 17.79 | 25.23 | |
| Actionformer+SADA [21] | 30.54 | 29.31 | 27.36 | 23.75 | 18.80 | 25.95 | |
| Tridet+SADA [10] | 31.34 | 30.16 | 27.94 | 24.48 | 19.31 | 26.64 | |
@@
| Ours (source-only) | 30.74 | 29.47 | 27.23 | 23.53 | 18.07 | 25.80 | |
| Actionformer+SADA [21] | 31.71 | 30.30 | 27.95 | 24.08 | 18.69 | 26.55 | |
| Tridet+SADA [10] | 29.55 | 28.27 | 26.16 | 22.89 | 17.63 | 24.90 | |
@@
| Ours (source-only) | 34.41 | 33.38 | 30.58 | 26.99 | 21.00 | 29.27 | |
| Actionformer+SADA [21] | 34.11 | 32.87 | 30.60 | 27.05 | 20.93 | 29.11 | |
| Tridet+SADA [10] | 34.50 | 33.19 | 30.65 | 27.25 | 20.83 | 29.29 | |
@@
| Ours (source-only) | 29.65 | 28.69 | 26.86 | 23.88 | 19.14 | 25.64 | |
| Actionformer+SADA [21] | 29.84 | 28.89 | 26.83 | 23.76 | 18.85 | 25.63 | |
| Tridet+SADA [10] | 30.21 | 29.33 | 27.58 | 24.35 | 19.44 | 26.18 | |
| Ours (SADA) | 31.54 | 30.68 | 28.77 | 25.52 | 20.22 | 27.34 | |
@@
| Ours (source-only) | 25.58 | 24.79 | 23.08 | 19.56 | 15.15 | 21.63 | |
| Actionformer+SADA [21] | 24.85 | 23.99 | 22.21 | 19.11 | 15.28 | 21.09 | |
| Tridet+SADA [10] | 24.88 | 24.00 | 22.20 | 19.02 | 14.88 | 21.00 | |
@@
| Ours (source-only) | 20.96 | 20.22 | 19.08 | 16.97 | 14.09 | 18.27 | |
| Actionformer+SADA [21] | 23.05 | 22.10 | 20.71 | 18.31 | 14.72 | 19.78 | |
| Tridet+SADA [10] | 21.00 | 20.08 | 18.64 | 16.39 | 13.13 | 17.85 | |
@@
In this section, we present the main experimental results evaluating appearance and acquisition shifts. All these experiments follow the standard transductive unsupervised DA protocol [63, 64], and report the mean average precision (mAP) at different intersections over union (IOU) thresholds ().
EpicKitchens100. In Table I we first present the results obtained in the 4 different scenarios that we designed to evaluate the performance of our method when facing different appearance shifts induced by changes in the background information. Concretely, we first evaluate the performance of the Actionformer [21] and TriDet [10], the two best-performing methods on EpicKitchens100 dataset [14], as well as our proposed architecture without our SADA loss. We then compare these 3 architectures with their extensions which include the SADA loss. Observe that our proposed loss improves the source-only version of all the architectures in all these 4 scenarios, improving by up to a mAP when evaluating our architecture on S1. Additionally, we observe that our final architecture –i.e., Ours(SADA)– yields an improvement over the second-best performing source-only method of up to a mAP for the black-counter kitchens scenarios and of mAP for the white-counter kitchens ones.
Similarly, the last two scenarios of Table I depict the setups containing acquisition domain shifts. These scenarios present a similar behavior than before. In all but one case on the Tridet [10] architecture, the use of the SADA loss yields a performance gain of up to mAP. This is a relative improvement. Our final model, moreover, improves the second-best performing source-only baselines by mAP and mAP, respectively.
| Model | mAP10% | mAP20% | mAP30% | mAP40% | mAP50% | Avg |
| Actionformer [21] | 31.22 | 28.51 | 23.82 | 18.91 | 13.94 | 23.28 |
| Tridet [10] | 30.18 | 27.06 | 22.58 | 17.81 | 13.10 | 22.15 |
| Ours (source-only) | 30.68 | 27.74 | 23.41 | 18.46 | 13.58 | 22.77 |
| Actionformer+SADA [21] | 31.46 | 28.57 | 24.17 | 19.04 | 13.87 | 23.42 |
| Tridet+SADA [10] | 31.53 | 28.41 | 24.04 | 18.88 | 13.98 | 23.37 |
| Ours (SADA) | 31.68 | 28.64 | 24.09 | 19.06 | 14.14 | 23.52 |
CharadesEgo. In Table II we report a similar experimental comparison of our proposed loss when evaluated on CharadesEgo. Concretely, we showcase the performance boost that SADA reports on the three test architectures –i.e., Actionformer [21], Tridet [10] and Ours. In the case of Tridet [10], for instance, the improvement reaches a mAP. In our proposed architecture, SADA yields an improvement over the source-only version of mAP. We additionally highlight that our proposed architecture Ours (SADA) attains the state-of-the-art results over all the baseline methods.
@@ 4.3 Ablation studies
Comparing to other domain adaptation methods. In Sec. 4.2 we showed that SADA consistently improves the performance of the three tested state-of-the-art source-only architectures when evaluated on our newly proposed setups. The question remains however of how well does SADA perform compared to existing domain adaptation methods. In this regard, following existing video-based domain adaptation works, we first evaluate various canonical UDA domain adaptation methods –i.e., ADDA [40], WDGRL [42], DANN [1] and MSTN [2]. These methods are integrated into our proposed underlying architecture resulting in a fair comparison with SADA. Additionally, we find that to the best of our knowledge, there is no existing UDA method for TAL in the literature that is directly comparable to this work. Nevertheless, to provide a richer comparison, we adapt the closest action segmentation proposal, SSTDA [7] and a state-of-the-art domain-adaptation method for video classification –i.e., TranSVAE [61]– to our proposed setup. We refer to the Supp. for all the details of these baselines.
Concretely, in Table III we compare SADA with these baselines on S1 and S2. These results empirically demonstrate the effectiveness of our method by attaining the best results over all tested UDA methods. More in detail, SADA yields an improvement of up to mAP and mAP on each of the scenarios respectively, and a mAP and mAP over the second best-performing method.
@@
| MSTN [2] | 31.07 | 27.16 | 17.21 | 25.14 |
| SSTDA [7] | 31.17 | 28.01 | 18.98 | 26.05 |
| TranSVAE [65] | 29.91 | 26.12 | 16.16 | 24.06 |
@@
| MSTN [2] | 30.88 | 27.44 | 18.25 | 25.52 |
| SSTDA [7] | 29.92 | 26.49 | 17.55 | 24.65 |
| TranSVAE [65] | 25.88 | 21.74 | 13.12 | 20.25 |
@@
Analysis of our loss. Next, we ablate over different variants of our proposed SADA loss (see Eq. 9). Concretely, in Table IV we study the effect of class-wise distribution alignment (Eq. 7), global distribution alignment (equivalent to DANN [1]) and semantic background alignment (Eq. 8). In this regard, we highlight that the global adaptation consistently improves upon aligning only background embeddings. This is because the latter yields only a partial alignment of the embeddings, not considering class anchors. Therefore, a rougher yet complete adaptation might seem beneficial. We also observe that aligning local class-wise distributions has a considerable positive effect, especially in the first scenario. Its effect, however, considerably decreases when combined with a global alignment [1], as all the class anchors are then subject to the concurrent alignments of domain-level and class-wise distributions. This observation is also consistent with the performance decrease that we observe when combining global and background alignment, which again suggests that the concurrent alignment of background embeddings with two adaptation losses is harmful to performance. Finally, we observe that we consistently obtain the best results when the local alignment loss is coupled with the complementary (non-overlapping) background loss – i.e., SADA loss – indicating that this semantic fine-grained, yet complete, alignment is the most desirable approach.
@@
Ablation of the effect of the hyper parameters. As described in Eq. 9, we define our SADA loss using a set of hyperparameters , where each of the parameters controls the influence of a given resolution level in the overall loss. Given the relevance of these parameters, below we ablate over different variations to showcase the effect that they have in the final performance of the model.
Concretely, we define the three following scenarios:
1.
All levels to 1: This scenario sets all the parameters , keeping the contribution of each of the levels the same.
2.
First 3 levels 1: This scenario sets the first half of the levels to 1 and the rest to 0. This is and .
3.
Last 3 levels 1: This scenario sets the first half of the levels to 0 and the rest to 1. This is and .
All in all, this study attempts to clarify how sensitive is our model to this design decision. In this regard, these results (see Table V) indicate a mild sensitivity to this choice. Specifically, we observe that the second best-performing method attains an absolute difference over the optimal choice of only mAP and mAP for the two considered scenarios S1 and S3, respectively. Moreover, we observe that the worst performing naive strategy still outperforms in both scenarios 4 out of the 6 tested UDA baselines.
[…]
| Strategy | mAP {10,30,50}% | Avg | ||
| All levs 1 | 32.42 | 29.59 | 19.34 | 27.11 |
| First 3 levs 1 | 30.96 | 27.48 | 18.29 | 25.58 |
| Last 3 levs 1 | 30.54 | 27.19 | 18.28 | 25.34 |
| ”Optimal choice” | 32.69 | 29.17 | 19.72 | 27.19 |
[…]
| Strategy | mAP {10,30,50}% | Avg | ||
| All levs 1 | 30.60 | 27.36 | 18.41 | 25.45 |
| First 3 levs 1 | 31.05 | 27.64 | 18.30 | 25.66 |
| Last 3 levs 1 | 30.95 | 27.59 | 18.56 | 25.70 |
| ”Optimal choice” | 31.60 | 28.22 | 18.98 | 26.27 |
Study of the class embedding. One of the main contributions of our work is to adversarially align distributions in a class-wise fashion. Intuitively, this requires that our level-wise domain discriminator knows the class distribution that it is aligning. In this regard, in Sec. 3.5 we propose to concatenate to every anchor a learnable embeddings of its corresponding class (see Eq. 7). In this ablation, we empirically justify our choice. For this, we compare our approach against several non-learnable alternatives to encode a given class : a naive one-hot encoding, a random class-wise dense initialization, and the sinusoidal encoding originally proposed by [33].
In Table VI we show the experimental results of each of the variants tested in S1 and S3. These indicate that a naive one-hot encoding of a class is the best-performing non-learnable strategy, obtaining considerable improvements over the other two tested non-learnable baselines. This improvement is especially prominent in the first scenario, boosting the performance over the other non-learnable strategies by mAP and mAP, respectively. Moreover, to our surprise, the popular Sinusoidal encoding [33] proves to be the worst-performing method, being consistently outperformed by even the random dense encoding. Finally, we highlight that in both scenarios the use of a learnable class embedding yields the best results, improving by mAP and mAP the performance of the one-hot encoding strategy, which justifies our choice.